OpenStudy (anonymous):

Is there a way to get bigger parentheses? For when I use fractions, the parentheses look a bit strange since they're shorter than the fractions.

OpenStudy (kinggeorge):

Use \left( and \right)\[\left(\frac{a}{b}\right)\]This also works with the other brackets available.

myininaya (myininaya):

\[\left( \frac{a}{b} \right)\] neat! :)

OpenStudy (anonymous):

\[\left(\frac ab\right)\]Yup. It works alright :)

myininaya (myininaya):

\[\left[ \frac{c+d}{a+b} \right]\]

OpenStudy (kinggeorge):

This is also a good way to use brackets on matrices. \[\left[\begin{matrix}1&2&3&4\\2&3&4&5\\3&4&5&6\\4&5&6&7\end{matrix}\right]\]

OpenStudy (kinggeorge):

You can also use \pmatrix and \bmatrix for matrices though.[\begin{pmatrix}1&2&3\\2&3&4\\3&4&5 \end{pmatrix}\[\begin{bmatrix}1&2&3\\2&3&4\\3&4&5 \end{bmatrix}\]

OpenStudy (anonymous):

\[\left(\frac ab\right)^{2^{3^{4^{5}}}}\]Exponents work too :) I obviously have a lot of \(\LaTeX\) to learn :)